# Reducing leakage of single-qubit gates for superconducting quantum processors using analytical control pulse envelopes

### v1

@@ II.3 Control pulses based on higher-derivative (HD) extension of DRAG
In this section, we present an alternative pulse shaping method for implementing fast, low-leakage single-qubit gates by extending the theoretical ideas of Ref. [14], in which a higher-derivative version of the DRAG method was theoretically proposed to mitigate multiple leakage transitions. In contrast to Ref. [14], we obtain $\Omega_{\mathrm{Q}}(t)$ from $\Omega_{\mathrm{I}}(t)$ using standard DRAG, decoupling the parameters of the in-phase and quadrature components, thus simplifying the calibration. We then use the higher-order derivatives to achieve a strong spectral suppression at the $ef$-transition instead of suppressing multiple separate transitions. Furthermore, we propose an ansatz based on a simple cosine series to ensure the continuity of the obtained control envelopes.
@@
Figure 2: Higher-derivative (HD) DRAG. (a) In-phase ($I$, solid) and quadrature ($Q$, dashed) envelopes of 6-ns HD DRAG pulses based on different maximum order of derivatives sampled (square markers) for our AWG and compared to a pulse without DRAG. The pulse without DRAG sets $\Omega_{\mathrm{Q}}=0$, the 1st-derivative DRAG pulse equals the conventional DRAG, the 2nd-derivative DRAG pulse sets $\Omega_{\mathrm{I}}$ according to Eq. (6) with $\Omega_{\mathrm{Q}}=0$, and the 3rd-derivative DRAG pulse is based on Eqs. (6) and (7). (b) Amplitude spectrum $|\hat{\Omega}_{\mathrm{IQ}}(f)|$ of the complex $IQ$ envelope $\Omega_{\mathrm{IQ}}$ for the same pulses, with the qubit anharmonicity $\alpha/(2\pi)$ suppressed by DRAG shown with the dashed vertical line and the inset presenting a close-up around $\alpha/(2\pi)$.
@@
Importantly, the basis function $g(t)$ and its three first derivatives should be continuous to ensure smooth control pulses, which we achieve by proposing the use of a short cosine series given by $g(t)=1-4/3\cos(2\pi t/t_{\mathrm{p}})+1/3\cos(4\pi t/t_{\mathrm{p}})$. The resulting envelopes of a third-derivative HD DRAG pulse in comparison to lower-derivative versions of DRAG are shown in Fig. 2(a). Due to expressing $g(t)$ as a cosine series, this version of HD DRAG can be viewed as an alternative way to obtain the coefficients $\{c_{n}\}$ of Eq. (2). As HD DRAG shares similar design criteria with FAST DRAG, the methods have many common advantages, such as the ability to implement strong spectral suppression around leakage transitions as illustrated in Fig. 2(b). However, HD DRAG has some downsides compared to FAST DRAG, such as inability to control the pulse bandwidth, the average power, or the exact strength and width of the spectral suppression.
@@ III Experimental implementation of fast, low-leakage single-qubit gates
To experimentally demonstrate the proposed pulse shaping methods, we benchmark the error and leakage per gate for fast single-qubit gates implemented using FAST DRAG and HD DRAG and compare the results to conventional pulse shapes, such as Cosine DRAG. In this work, we utilize a two-qubit test device with a design and fabrication similar to Ref. [4]. We focus on one of the two flux-tunable qubits that is operated at its sweet-spot with a frequency of 4.417 GHz, an anharmonicity of -212 MHz, and average coherence times of about 37 us and 22 us for $T_{1}$ and $T_{2}^{\mathrm{e}}$ (Hahn echo), respectively. We implement single-qubit rotations by sending microwave pulses to a charge line coupled to the qubit through a targeted capacitance of 0.12 fF. The microwave pulses are generated by up-converting intermediate-frequency pulses from a Zurich Instruments HDAWG using an IQ-mixer, and then propagated to the sample attached to the cold-plate of a dilution refrigerator (Bluefors LD400) through coaxial lines with 67 dB room-temperature attenuation at 4.5 GHz, see Fig. 6 in Appendix D. In this work, we focus on the calibration of $R_{X}(\pi/2)$-gates to avoid issues caused by drive non-linearities [38]. Furthermore, $R_{X}(\pi/2)$-gates along with Virtual-$Z$ rotations conveniently enable the creation of arbitrary-angle single-qubit gates with at most two $R_{X}(\pi/2)$ gates [16].
@@
To implement single-qubit gates using FAST DRAG, we suppress the spectrum of the in-phase component across a baseband frequency interval of $[f_{\mathrm{l},1},f_{\mathrm{h},1}]=[f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}}]=[194,214]$ MHz covering the qubit anharmonicity, while using another frequency interval of $[f_{\mathrm{l},2},f_{\mathrm{h},2}]=[450,1000]$ MHz effectively setting a cutoff frequency of $f_{\mathrm{c}}=450$ MHz. In addition, we use a weight ratio of $w_{\mathrm{ef}}=w_{1}/w_{2}=5$ ($w_{\mathrm{ef}}=100$) and set the number of cosine terms to $N=4$ ($N=5$) for FAST DRAG-L (DRAG-P). The parameter values are chosen based on prior simulations to provide a low leakage rate across the studied gate durations $t_{\mathrm{g}}\in[5,20]$ ns, and a higher weight ratio is used for FAST DRAG-P to enable sufficient spectral suppression even though the DRAG coefficient is optimized for correcting phase errors. As detailed in Appendix I, we sweep the parameters of FAST DRAG pulses to demonstrate that the leakage error stays low under moderate changes to the parameter values. For HD DRAG, we set $\beta_{2}=1/\alpha^{2}$ to minimize the spectral energy density of the in-phase envelope at the anharmonicity. In all of the results of this section, we include a 0.41-ns delay between consecutive pulses as a part of the reported gate duration $t_{\mathrm{g}}$, i.e., $t_{\mathrm{g}}=t_{\mathrm{p}}+0.41$ ns, and apply predistortion for the control pulse envelopes as discussed in more detail in Sec. III.1 and Appendix G.
@@
To compare the error and leakage per gate for the proposed and conventional pulse shapes, we calibrate the single-qubit gate parameters $A$, $\beta$, $\omega_{\mathrm{d}}/(2\pi)$ and the virtual Z-rotation phase increment $\varphi_{\mathrm{z}}$ as explained in Appendix F, and benchmark the resulting gate performance using randomized benchmarking (RB) [39, 40, 41, 42] combined with 3-state discrimination (see Appendix E) to further enable leakage RB analysis [15, 43]. The RB protocol enables estimating the average gate error by fitting an exponential model $p_{\mathrm{g}}(N_{\mathrm{Cl}})=A_{\mathrm{g}}+B_{\mathrm{g}}p^{N_{\mathrm{Cl}}}$ to the ground state probability $p_{\mathrm{g}}$ measured as a function of the Clifford sequence length $N_{\mathrm{Cl}}$ and then extracting the average error per Clifford as $\varepsilon_{\mathrm{Cl}}=(1-p)/2$. Since we decompose the Clifford gates using on average $N_{\mathrm{g}}\approx 2.21$ native gates from the gate set $\{I,R_{X}(\pm\pi/2),R_{Y}(\pm\pi/2)\}$ [42], we estimate the error per gate as $\varepsilon_{\mathrm{g}}=\varepsilon_{\mathrm{Cl}}/N_{\mathrm{g}}$. We further estimate the leakage error by fitting an exponential model $p_{\mathrm{f}}(N_{\mathrm{Cl}})=A_{\mathrm{f}}+B_{\mathrm{f}}\lambda_{1}^{N_{\mathrm{Cl}}}$ to the $f$-state population and then computing the leakage per gate as $L_{\mathrm{g}}=A_{\mathrm{f}}(1-\lambda_{1})/N_{\mathrm{g}}$.
@@
From the measured leakage error shown in Fig. 3(c), we confirm that the error for short gates is dominated by leakage. We also find a good agreement between the measured leakage error and Master equation simulations based on a 4-level Duffing oscillator model even though the simulations ignore any potential non-idealities of the control electronics, such as finite bandwidth, dynamic range or sampling rate, see Appendix C for more details. Importantly, the experimental leakage per gate stays at or below $3.0\times 10^{-5}$ down to a gate duration of 6.25 ns for both FAST DRAG-L and HD DRAG-L, providing up to a 20-fold reduction in leakage compared to Cosine DRAG-L. As shown in the example RB experiment of Figs. 3(d) and (e), FAST DRAG-L with a gate duration of $t_{\mathrm{g}}=6.25$ ns reduces the error and leakage per gate by factors of 4 and 23, respectively, compared to Cosine DRAG-L with equivalent duration. The average leakage per gate of $(2.7\pm 0.3)\times 10^{-5}$ and the corresponding leakage per Clifford of $(6\pm 1)\times 10^{-5}$ obtained using FAST DRAG-L at $t_{\mathrm{g}}=6.25$ ns, i.e., $t_{\mathrm{g}}\alpha/(2\pi)=1.33$, are lower compared to the lowest value we have found in literature for such short pulses [26]. In Ref. [26], closed-loop optimization of a piecewise constant $R_{X}(\pi/2)$-pulse resulted in an order of magnitude higher leakage per Clifford of $4.4\times 10^{-4}$ using a transmon with an anharmonity of $-315$ MHz and a gate duration of 4.16 ns corresponding to a normalized gate duration of $t_{\mathrm{g}}\alpha/(2\pi)=1.31$ practically equivalent to our work.
@@
To compare the gate performance of the novel pulse shapes against other conventional pulses beyond cosine DRAG, we further experimentally benchmark the leakage rate for DRAG pulses based on Gaussian and Slepian in-phase envelopes. For the Gaussian envelope, we use $t_{\mathrm{p}}=5\sigma$ ns and subtract the discontinuity similar to, e.g., Ref. [22]. To construct the Slepian envelope, we use FAST pulse shaping with an effective cutoff frequency of $185~{}\mathrm{MHz}<\alpha/(2\pi)$. Further results obtained for Gaussian and Slepian envelopes with different parameter choices are provided in Appendices H and I. Based on measurements of leakage per gate as a function of gate duration in Appendix H, we estimate an experimental speed limit for each of the control pulses by determining the shortest gate duration with a leakage per gate below $5\times 10^{-5}$. As shown in Fig. 4(a), FAST DRAG-L and HD DRAG-L enable the fastest low-leakage gates, reaching the leakage threshold at a gate duration of approximately 6 ns. The corresponding speed limits are 7.5 ns, 8.7 ns, and 10.4 ns for Slepian DRAG-L, Cosine DRAG-L, and Gaussian DRAG-L. When calibrating DRAG to minimize phase errors, the corresponding speed limit is 12-13 ns for all the other pulses than FAST DRAG-P, for which the leakage rate stays below the threshold down to a gate duration of 9.6 ns thanks to a strong spectral suppression around the $ef$-transition provided by FAST pulse shaping.
@@ III.1 Characterization and pre-compensation of control pulse distortions
As mentioned in the previous section, coherent errors related to microwave pulse distortions appear to prevent us from implementing coherence-limited sub-10-ns gates. Here, we discuss experimental evidence of such pulse distortions using the qubit as a sensor and present our approach for predistorting control pulse envelopes to improve gate fidelity. We classify the pulse distortions into two types based on their effect on the pulse envelopes similarly to Ref. [28], see Fig. 5(a). The first type of distortion, named intra-quadrature distortion (I-distortion), corresponds to a distortion of the in-phase (quadrature) envelope resulting from a pulse in the in-phase (quadrature) envelope. The second type of distortion, named cross-quadrature distortion (C-distortion), refers to a distortion of the quadrature (in-phase) envelope caused by a pulse in the in-phase (quadrature) envelope. Both types of distortions result in coherent errors, the effects of which depend on previous gates and the delay $t_{\mathrm{d}}$ between pulses in a way that cannot be corrected by optimizing the sample amplitudes during the gate, e.g., as in Ref. [26]. As such errors are non-Markovian, the gates cannot be described by fixed Pauli transfer matrices and thus, the errors cannot be characterized using, e.g., gate set tomography [46, 47]. However, it is possible to design gate sequences amplifying errors related to each type of distortion such that the gate sequences are almost exclusively sensitive to only one type of distortion.
@@
To characterize C-distortions, we use a circuit consisting of repeated ($R_{X}(\pi),R_{X}(-\pi)$)-pairs as suggested in Ref. [28] and append the sequence with a final $R_{Y}(\pi/2)$-gate to make the sequence linearly sensitive to phase errors, see Fig. 5(a). As a result, the sequence is efficient in detecting C-distortions causing phase errors, the magnitude of which varies depending on the delay $t_{\mathrm{d}}$ between pulses. For the characterization of I-distortions, we propose a circuit consisting of repeated ($R_{X}(\pi),R_{Y+\varphi}(\pi)$)-pairs prepended with a $R_{X}(\pi/2)$ gate and appended with a $R_{Y}(\pi/2)$ gate to make the measured excited state probability linearly sensitive to distortion-induced coherent errors. In this experiment, each pair of ($R_{X}(\pi),R_{Y}(\pi)$)-gates would ideally result in a $Z$-rotation. In the presence of I-distortions, the tail of the previous pulse tilts the rotation axis of the following pulse, resulting in a change of the relative angle between the axes of $R_{X}(\pi)$ and $R_{Y}(\pi)$ as shown in Fig 5(b). This further causes the resulting $Z$-rotation to have an over-rotation that can be amplified with repeated gates. Other common types of errors, such as amplitude miscalibration or phase errors, do not result in such overrotation to first order, but instead change the resulting rotation axis and thus do not contribute to the error amplification. As shown in Fig 5(b), we sweep the relative phase between $R_{X}(\pi)$ and $R_{Y+\varphi}(\pi)$ through $\varphi$ in order to determine an axis shift angle $\varphi_{\mathrm{s}}$ cancelling the axis tilt caused by the pulse tails. Thus, we regard the axis shift $\varphi_{\mathrm{s}}$ as a measure of the magnitude of the I-distortions. Furthermore, we accompany the I- and C-distortion characterization experiments with a sweep of the delay between pulses to characterize the temporal profile of the pulse distortions.
@@
We mitigate I-distortions by separately predistorting the $I$ and $Q$ control pulse envelopes using an exponential infinite impulse response (IIR) inverse filter in analogy to the predistortion of flux pulses [48, 49, 5], see Fig. 5(c) and Appendix G for more details. This choice of the filter function is motivated by the observation that the axis shift $\varphi_{\mathrm{s}}$ measured from the I-distortion characterization decays approximately exponentially as a function of the delay between pulses as shown in Fig. 5(d). Using calibration measurements explained in Appendix G, we obtain a time constant $\tau_{1}=8$ ns and amplitude coefficient $a_{1}=-0.028$ as the predistortion parameters used for both $I$ and $Q$ envelopes. When using predistortion, we observe that the exponential dependence is removed in the I-distortion characterization resulting in a significantly smaller axis shift as shown in Fig. 5(d), providing support for the used predistortion model. In Fig 5(e), we further show that the predistortion significantly reduces the axis shift $\varphi_{\mathrm{s}}$ for FAST DRAG-P and FAST DRAG-L pulses across gate durations $t_{\mathrm{g}}\in[6,20]$ ns. For FAST DRAG-L, the remaining axis shift is higher since the virtual-$Z$ rotations preceding each pulse effectively convert part of the non-corrected C-distortion into I-distortion. In Fig. 5(f), we compare the RB gate error obtained using FAST DRAG-L pulses with and without predistortion and demonstrate that the gate error is systematically reduced by $(2-6)\times 10^{-5}$ across the tested gate durations between 6 ns and 11 ns.
@@
When applying the predistortion, we still find a small coherent error as discussed above and illustrated in Fig. 4(c). We attribute this remaining coherent error to C-distortions that we do not correct in this work. To verify this conclusion, we conduct a C-distortion characterization experiment for varying pulse-to-pulse delays using a FAST DRAG-P pulse with $t_{\mathrm{p}}=9.17$ ns. As shown in Fig. 5(g), the gates indeed suffer from delay-dependent phase errors that we attribute to remaining C-distortions. Note that we do not use the measurement scheme proposed in Ref. [28] to obtain the finite impulse response filter corresponding to the C-distortions since the assumption of vanishing gate duration does not hold well for a transmon-based system unlike for the flux qubit studied in Ref. [28].

### v2

@@ II.3 Control pulses based on higher-derivative (HD) extension of DRAG
In this section, we present an alternative pulse shaping method for implementing fast, low-leakage single-qubit gates by extending the theoretical ideas of Ref. [16], in which a higher-derivative version of the DRAG method was theoretically proposed to mitigate multiple leakage transitions. In contrast to Ref. [16], we obtain $\Omega_{\mathrm{Q}}(t)$ from $\Omega_{\mathrm{I}}(t)$ using standard DRAG, decoupling the parameters of the in-phase and quadrature components, thus simplifying the calibration. Furthermore, we use the higher-order derivatives to achieve a strong spectral suppression of the dominant leakage transition, i.e., the $ef$-transition instead of suppressing multiple separate transitions as in Ref. [16]. We also propose an ansatz based on a simple cosine series to ensure the continuity of the obtained control envelopes.
@@
Figure 2: Higher-derivative (HD) DRAG. (a) In-phase ($I$, solid) and quadrature ($Q$, dashed) envelopes of 6-ns HD DRAG pulses based on different maximum order of derivatives sampled (square markers) for our AWG and compared to a pulse without DRAG. The pulse without DRAG sets $\Omega_{\mathrm{Q}}=0$, the 1st-derivative DRAG pulse equals the conventional DRAG, the 2nd-derivative DRAG pulse sets $\Omega_{\mathrm{I}}$ according to Eq. (6) with $\Omega_{\mathrm{Q}}=0$ similarly to the 2nd-derivative DRAG solution of Ref. [16] , and the 3rd-derivative DRAG pulse is based on Eqs. (6) and (7). (b) Amplitude spectrum $|\hat{\Omega}_{\mathrm{IQ}}(f)|$ of the complex $IQ$ envelope $\Omega_{\mathrm{IQ}}$ for the same pulses, with the qubit anharmonicity $\alpha/(2\pi)$ suppressed by DRAG shown with the dashed vertical line and the inset presenting a close-up around $\alpha/(2\pi)$.
@@
Importantly, the basis function $g(t)$ and its three first derivatives should be continuous to ensure smooth control pulses, which we achieve by proposing the use of a short cosine series given by $g(t)=1-4/3\cos(2\pi t/t_{\mathrm{p}})+1/3\cos(4\pi t/t_{\mathrm{p}})$. The resulting envelopes of a third-derivative HD DRAG pulse in comparison to lower-derivative versions of DRAG are shown in Fig. 2(a). After expressing $g(t)$ as a cosine series, this version of HD DRAG is a special case of FAST DRAG with $N=2$ and a single suppressed frequency interval defined by $(f_{\mathrm{l},\mathrm{ef}}+f_{\mathrm{h},\mathrm{ef}})/2=|\alpha|/(2\pi)$ and $f_{\mathrm{h},\mathrm{ef}}-f_{\mathrm{l},\mathrm{ef}}\to 0$.
Thus, HD DRAG and FAST DRAG have many common advantages since they share similar design criteria. Most importantly, HD DRAG enables strong spectral suppression of leakage transitions similarly to FAST DRAG as illustrated in Fig. 2(b). Multiple leakage transitions can also be suppressed with HD DRAG by introducing higher-order derivatives as explained in Appendix B. Despite these similarities, there are also several key differences between the two methods. HD DRAG has less parameters than FAST DRAG, and the third-derivative HD DRAG studied in this work only introduces a single additional parameter $\beta_{2}$ determined by the measured anharmonicity $\alpha/(2\pi)$. However, HD DRAG does not allow straightforward control over the pulse bandwidth, the average power, or the exact strength and width of the spectral suppression unlike FAST DRAG, which offers more flexibility for the pulse design. Hence, we foresee that the flexibility of FAST DRAG is beneficial especially in scenarios with multiple leakage transitions, such as in simultaneous single-qubit gates as briefly discussed in Appendix F.
@@ III Experimental implementation of fast, low-leakage single-qubit gates
To experimentally demonstrate the proposed pulse shaping methods, we benchmark the error and leakage per gate for fast single-qubit gates implemented using FAST DRAG and HD DRAG and compare the results to conventional pulse shapes, such as Cosine DRAG. In this work, we utilize a two-qubit test device with a design and fabrication similar to Ref. [4]. We focus on one of the two flux-tunable qubits that is operated at its sweet-spot with a frequency of 4.417 GHz, an anharmonicity of -212 MHz, and average coherence times of about 37 us and 22 us for $T_{1}$ and $T_{2}^{\mathrm{e}}$ (Hahn echo), respectively. We implement single-qubit rotations by sending microwave pulses to a charge line coupled to the qubit through a targeted capacitance of 0.12 fF. The microwave pulses are generated by up-converting intermediate-frequency pulses from a Zurich Instruments HDAWG using an IQ-mixer, and then propagated to the sample attached to the cold-plate of a dilution refrigerator (Bluefors LD400) through coaxial lines with 67 dB room-temperature attenuation at 4.5 GHz, see Fig. 6 in Appendix D. Since our gate decomposition is based on $R_{X}(\pi/2)$-gates, it is sufficient to calibrate the pulse amplitude for a single rotation angle $\theta=\pi/2$ allowing us to neglect drive non-linearities [52].
@@
To implement single-qubit gates using FAST DRAG, we suppress the spectrum of the in-phase component across a baseband frequency interval of $[f_{\mathrm{l},1},f_{\mathrm{h},1}]=[f_{\mathrm{l},\mathrm{ef}},f_{\mathrm{h},\mathrm{ef}}]=[194,214]$ MHz covering the qubit anharmonicity, while using another frequency interval of $[f_{\mathrm{l},2},f_{\mathrm{h},2}]=[450,1000]$ MHz effectively setting a cutoff frequency of $f_{\mathrm{c}}=450$ MHz. In addition, we use a weight ratio of $w_{\mathrm{ef}}=w_{1}/w_{2}=5$ ($w_{\mathrm{ef}}=100$) and set the number of cosine terms to $N=4$ ($N=5$) for FAST DRAG-L (DRAG-P). Here, a higher weight ratio is used for FAST DRAG-P to enable sufficient spectral suppression as the DRAG coefficient is reserved for correcting phase errors. The parameter values are chosen using physics-based heuristics and prior simulations to provide a low leakage rate across the studied gate durations $t_{\mathrm{g}}\in[5,20]$ ns. See Appendix F for general heuristic guidelines on the selection of these hyperparameters and Appendix J for parameter sweeps demonstrating that the leakage error stays low under moderate changes to the hyperparameter values. For HD DRAG, we set $\beta_{2}=1/\alpha^{2}$ to minimize the spectral energy density of the in-phase envelope at the anharmonicity. In all of the results of this section, we include a 0.41-ns delay between consecutive pulses as a part of the reported gate duration $t_{\mathrm{g}}$, i.e., $t_{\mathrm{g}}=t_{\mathrm{p}}+0.41$ ns, and apply predistortion for the control pulse envelopes as discussed in more detail in Sec. III.1 and Appendix H.
@@
To compare the error and leakage per gate for the proposed and conventional pulse shapes, we calibrate the single-qubit gate parameters $A$, $\beta$, $\omega_{\mathrm{d}}/(2\pi)$ and the virtual Z-rotation phase increment $\varphi_{\mathrm{z}}$ using sequences of error amplification experiments without any iterative closed-loop optimization as explained in Appendix G, and benchmark the resulting gate performance using randomized benchmarking (RB) [53, 54, 55, 56] combined with 3-state discrimination (see Appendix E) to further enable leakage RB analysis [17, 57]. The RB protocol enables estimating the average gate error by fitting an exponential model $p_{\mathrm{g}}(N_{\mathrm{Cl}})=A_{\mathrm{g}}+B_{\mathrm{g}}p^{N_{\mathrm{Cl}}}$ to the ground state probability $p_{\mathrm{g}}$ measured as a function of the Clifford sequence length $N_{\mathrm{Cl}}$ and then extracting the average error per Clifford as $\varepsilon_{\mathrm{Cl}}=(1-p)/2$. Since we decompose the Clifford gates using on average $N_{\mathrm{g}}\approx 2.21$ native gates from the gate set $\{I,R_{X}(\pm\pi/2),R_{Y}(\pm\pi/2)\}$ [56], we estimate the error per gate as $\varepsilon_{\mathrm{g}}=\varepsilon_{\mathrm{Cl}}/N_{\mathrm{g}}$ similarly to prior work using an equivalent Clifford decomposition [7, 3, 18]. We further estimate the leakage error by fitting an exponential model $p_{\mathrm{f}}(N_{\mathrm{Cl}})=A_{\mathrm{f}}+B_{\mathrm{f}}\lambda_{1}^{N_{\mathrm{Cl}}}$ to the $f$-state population and then computing the leakage per gate as $L_{\mathrm{g}}=A_{\mathrm{f}}(1-\lambda_{1})/N_{\mathrm{g}}$.
@@
From the measured leakage error shown in Fig. 3(c), we confirm that the error for short gates is dominated by leakage. We also find a good agreement between the measured leakage error and Master equation simulations based on a 4-level Duffing oscillator model even though the simulations ignore any potential non-idealities of the control electronics, such as finite bandwidth, dynamic range or sampling rate, see Appendix C for more details. Importantly, the experimental leakage per gate stays at or below $3.0\times 10^{-5}$ down to a gate duration of 6.25 ns for both FAST DRAG-L and HD DRAG-L, providing up to a 20-fold reduction in leakage compared to Cosine DRAG-L. However, the leakage error for FAST DRAG-L showcases a shallow local maximum around $t_{\mathrm{g}}|\alpha|/(2\pi)\approx 2$ since our heuristics for the hyperparameter selection have been primarily designed to enable low leakage errors for fast gates $t_{\mathrm{g}}|\alpha|/(2\pi)\in[1,2]$. As shown in the example RB experiment of Figs. 3(d) and (e), FAST DRAG-L with a gate duration of $t_{\mathrm{g}}=6.25$ ns reduces the error and leakage per gate by factors of 4 and 23, respectively, compared to Cosine DRAG-L with equivalent duration. The average leakage per gate of $(2.7\pm 0.3)\times 10^{-5}$ and the corresponding leakage per Clifford of $(6\pm 1)\times 10^{-5}$ obtained using FAST DRAG-L at $t_{\mathrm{g}}=6.25$ ns, i.e., $t_{\mathrm{g}}|\alpha|/(2\pi)=1.33$, are lower compared to the lowest value we have found in literature for such short pulses [7]. In Ref. [7], closed-loop optimization of a piecewise constant $R_{X}(\pi/2)$-pulse resulted in an order of magnitude higher leakage per Clifford of $4.4\times 10^{-4}$ using a transmon with an anharmonity of $-315$ MHz and a gate duration of 4.16 ns corresponding to a normalized gate duration of $t_{\mathrm{g}}|\alpha|/(2\pi)=1.31$ practically equivalent to our work.
@@
To compare the gate performance of the novel pulse shapes against other conventional pulses beyond cosine DRAG, we further experimentally benchmark the leakage rate for DRAG pulses based on Gaussian and Slepian in-phase envelopes. For the Gaussian envelope, we use $t_{\mathrm{p}}=5\sigma$ ns and subtract the discontinuity similar to, e.g., Ref. [24]. To construct the Slepian envelope, we use FAST pulse shaping with an effective cutoff frequency of $185~{}\mathrm{MHz}<\alpha/(2\pi)$. Further results obtained for Gaussian and Slepian envelopes with different parameter choices are provided in Appendices I and J. Based on measurements of leakage per gate as a function of gate duration in Appendix I, we estimate an experimental speed limit for each of the control pulses by determining the shortest gate duration with a leakage per gate below $5\times 10^{-5}$. As shown in Fig. 4(a), FAST DRAG-L and HD DRAG-L enable the fastest low-leakage gates, reaching the leakage threshold at a gate duration of approximately 6 ns. The corresponding speed limits are 7.5 ns, 8.7 ns, and 10.4 ns for Slepian DRAG-L, Cosine DRAG-L, and Gaussian DRAG-L. When calibrating DRAG to minimize phase errors, the corresponding speed limit is 12-13 ns for all the other pulses than FAST DRAG-P, for which the leakage rate stays below the threshold down to a gate duration of 9.6 ns thanks to a strong spectral suppression around the $ef$-transition provided by FAST pulse shaping.
@@ III.1 Characterization and pre-compensation of control pulse distortions
As mentioned in the previous section, coherent errors related to microwave pulse distortions appear to prevent us from implementing coherence-limited sub-10-ns gates. Here, we discuss experimental evidence of such pulse distortions using the qubit as a sensor and present our approach for predistorting control pulse envelopes to improve gate fidelity. We classify the pulse distortions into two types based on their effect on the pulse envelopes similarly to Ref. [45], see Fig. 5(a). The first type of distortion, named intra-quadrature distortion (I-distortion), corresponds to a distortion of the in-phase (quadrature) envelope resulting from a pulse in the in-phase (quadrature) envelope. The second type of distortion, named cross-quadrature distortion (C-distortion), refers to a distortion of the quadrature (in-phase) envelope caused by a pulse in the in-phase (quadrature) envelope. Both types of distortions result in coherent errors, the effects of which depend on previous gates and the delay $t_{\mathrm{d}}$ between pulses in a way that cannot be corrected by optimizing the sample amplitudes during the gate, e.g., as in Ref. [7]. As such errors are non-Markovian, the gates cannot be described by fixed Pauli transfer matrices extracted, e.g., from gate set tomography [60, 61], whereas the characterization of the full process tensor [62, 63, 64] capturing non-Markovian errors would be time consuming. Here, we instead characterize the I- and C-distortions using straightforward error amplification experiments which are almost exclusively sensitive to only one type of distortion.
@@
To characterize C-distortions, we use a circuit consisting of repeated ($R_{X}(\pi),R_{X}(-\pi)$)-pairs as suggested in Ref. [45] and append the sequence with a final $R_{Y}(\pi/2)$-gate to make the sequence linearly sensitive to phase errors, see Fig. 5(a). As a result, the sequence is efficient in detecting C-distortions causing phase errors, the magnitude of which varies depending on the delay $t_{\mathrm{d}}$ between pulses. For the characterization of I-distortions, we propose a circuit consisting of repeated ($R_{X}(\pi),R_{Y+\varphi}(\pi)$)-pairs prepended with a $R_{X}(\pi/2)$ gate and appended with a $R_{Y}(\pi/2)$ gate to make the measured excited state probability linearly sensitive to distortion-induced coherent errors. In this experiment, each pair of ($R_{X}(\pi),R_{Y}(\pi)$)-gates would ideally result in a $Z$-rotation. In the presence of I-distortions, the tail of the previous pulse tilts the rotation axis of the following pulse, resulting in a change of the relative angle between the axes of $R_{X}(\pi)$ and $R_{Y}(\pi)$, as shown in Fig 5(b). This further causes the resulting $Z$-rotation to have an over-rotation that can be amplified with repeated gates. Other common types of errors, such as amplitude miscalibration or phase errors, do not result in such overrotation to first order, but instead change the resulting rotation axis and thus do not contribute to the error amplification. We sweep the relative phase between $R_{X}(\pi)$ and $R_{Y+\varphi}(\pi)$ through $\varphi$ in order to determine an axis shift angle $\varphi_{\mathrm{s}}$ leading to a vanishing overrotation angle of the composite $Z$-rotation, thus indicating cancellation of the axis tilt caused by the pulse tails as schematically illustrated in Fig 5(b). Therefore, we regard the axis shift $\varphi_{\mathrm{s}}$ as a measure of the magnitude of the I-distortions. Furthermore, we accompany the I- and C-distortion characterization experiments with a sweep of the delay between pulses to characterize the temporal profile of the pulse distortions.
@@
We mitigate I-distortions by separately predistorting the $I$ and $Q$ control pulse envelopes using an exponential infinite impulse response (IIR) inverse filter in analogy to the predistortion of flux pulses [65, 66, 5], see Fig. 5(c) and Appendix H for more details. This choice of the filter function is motivated by the observation that the axis shift $\varphi_{\mathrm{s}}$ measured from the I-distortion characterization decays approximately exponentially as a function of the delay between pulses, as shown in Fig. 5(d). Using calibration measurements explained in Appendix H, we obtain a time constant $\tau_{1}=8$ ns and amplitude coefficient $a_{1}=-0.028$ as the predistortion parameters used for both $I$ and $Q$ envelopes. When using predistortion, we observe that the exponential dependence is removed in the I-distortion characterization resulting in a significantly smaller axis shift as shown in Fig. 5(d), providing support for the used predistortion model. In Fig 5(e), we further show that the predistortion significantly reduces the axis shift $\varphi_{\mathrm{s}}$ for FAST DRAG-P and FAST DRAG-L pulses across gate durations $t_{\mathrm{g}}\in[6,20]$ ns. For FAST DRAG-L, the remaining axis shift is higher since the virtual Z-rotations preceding each pulse effectively convert part of the non-corrected C-distortion into I-distortion. In Fig. 5(f), we compare the RB gate error obtained using FAST DRAG-L pulses with and without predistortion and demonstrate that the gate error is systematically reduced by $(2-6)\times 10^{-5}$ across the tested gate durations between 6 ns and 11 ns.
@@
When applying the predistortion, we still find a small coherent error as discussed above and illustrated in Fig. 4(c). We attribute this remaining coherent error to C-distortions that we were not able to correct in this work, see Appendix H. To verify this conclusion, we conduct a C-distortion characterization experiment for varying pulse-to-pulse delays using a FAST DRAG-P pulse with $t_{\mathrm{p}}=9.17$ ns. As shown in Fig. 5(g), the gates indeed suffer from delay-dependent phase errors that we attribute to remaining C-distortions. Note that we do not use the measurement scheme proposed in Ref. [45] to obtain the finite impulse response filter corresponding to the C-distortions since the assumption of vanishing gate duration does not hold well for a transmon-based system unlike for the flux qubit studied in Ref. [45].